When 50.0 mL of .10 M HCl and 50.0 mL of .10 M NaOH, both at 22 oC, are added to a calorimeter, the temperature of the mixture reaches 28.9 oC. Calculate the heat produced by this reaction.
Density of water 1.00g/mL. Specific heat of water = 4.184 J/g oC
Answer #1

Volume of solution = 50.0+50.0 = 100.0ml

Since density is given 1.00g/ml , mass of solution = 100.0g

∆T = 28.9-22 = 6.9℃

Heat produced = m.s.∆T = 100.0g × 4.184J/g.℃ × 6.9℃ = 2886.96 J